# Age-Aware Edge-Blind Federated Learning via Over-the-Air Aggregation

### v1

@@ II-B Age-Aware Over-the-Air Communication Model
$\bm{\theta}(t{+}1)=\bm{\theta}(t)+\widehat{\Delta\bm{\theta}}(t)$ (17)
@@ III-A Preliminaries
$\Gamma\triangleq F^{\ast}-\sum_{m=1}^{M}\frac{B_{m}}{B}\,F_{m}^{\ast}.$ (24)
@@ IV Experimental Results and Discussion
We first investigate the effect of the number of antennas $N$. We set $M=10$ clients and consider an extreme non-IID scenario in which each client is given training samples of only one MNIST data label. We set $\tau=3$ local SGD iterations to train and update the model. For compression, with a slight abuse of notation, the value of $r$ is set to $0.9$ of the total number of parameters (i.e., $r$ here represents a ratio); likewise, we set the value of $k$ to $0.2$ of the selected largest magnitudes. The wireless channel parameters are set to $\sigma_{h}^{2}=1$ and $\sigma_{z}^{2}=5$, while each client device is equipped with a power budget of $\bar{P}=10$. The results in Fig. 1 show that using only a single antenna is insufficient for the system to converge to an acceptable test accuracy. This is because the interference and noise effects of the received signals cannot be mitigated with only one antenna, leading to a significant degradation in training performance. In contrast, increasing the number of antenna elements significantly improves convergence speed and test accuracy, since having more antennas makes the reconstructed received signal at the PS closer to the true average of the client updates. The results in Fig. 1 also show that AgeTop-k achieves better convergence rate and accuracy than rTop-k by using relatively larger number of antennas ($N=50$ and $N=1000$), since this helps mitigate the randomness introduced by the channel, and represents relatively good channel conditions. However, when the number of the receiving antenna elements is relatively small ($N=10$), the two techniques achieve a similar performance, since the randomness introduced by the channel cannot be mitigated using only $10$ antenna elements.
@@ V Conclusion
This paper introduced an age-aware edge-blind over-the-air FL framework that is efficient and avoids the need for client-side CSI estimation. The PS aggregates the transmitted updates using multi-antenna MRC and updates only a small subset of coordinates through a two-step technique (AgeTop-$k$): selecting the $r$-largest model magnitudes, and then updating the $k$ most stale of which. By compressing each update into only $k$ selected coordinates that fit into a single OFDM symbol, the proposed method reduces latency while keeping the most useful information. Our convergence analysis has shown that increasing the number of PS antennas improves convergence rate and accuracy. The analysis has also demonstrated a trade-off between compression quality and transmission quality: larger $k$ reduces compression error but increases noise sensitivity, and hence an optimal value must be employed to balance both effects. Theoretical results have been validated by experimental results. Future work include analytically exploring the optimal value of $k$ based on channel conditions and compression error, as well as extending to a fully-blind (CSI-free) system at both the PS and clients.

### v2

@@ II-B Age-Aware Over-the-Air Communication Model
$\bm{\theta}(t{+}1)=\bm{\theta}(t)+\widehat{\Delta\bm{\theta}}(t).$ (17)
@@ III-A Preliminaries
$\Gamma\triangleq F^{\ast}-\sum_{m=1}^{M}\frac{B_{m}}{B}\,F_{m}^{\ast},$ (24)
@@ IV Experimental Results and Discussion
We first investigate the effect of the number of antennas $N$. We set $M=10$ clients and consider an extreme non-IID scenario in which each client is given training samples of only one MNIST data label. We set $\tau=3$ local SGD iterations to train and update the model. For compression, with a slight abuse of notation, the value of $r$ is set to $0.9$ of the total number of parameters (i.e., $r$ here represents a ratio); likewise, we set the value of $k$ to $0.2$ of the total number of parameters. The wireless channel parameters are set to $\sigma_{h}^{2}=1$ and $\sigma_{z}^{2}=5$, while each client device is equipped with a power budget of $\bar{P}=10$. The results in Fig. 1 show that using only a single antenna is insufficient for the system to converge to an acceptable test accuracy. This is because the interference and noise effects of the received signals cannot be mitigated with only one antenna, leading to a significant degradation in training performance. In contrast, increasing the number of antenna elements significantly improves convergence speed and test accuracy, since having more antennas makes the reconstructed received signal at the PS closer to the true average of the client updates. The results in Fig. 1 also show that AgeTop-k achieves better convergence rate and accuracy than rTop-k by using relatively larger number of antennas ($N=50$ and $N=1000$), since this helps mitigate the randomness introduced by the channel, and represents relatively good channel conditions. However, when the number of the receiving antenna elements is relatively small ($N=10$), the two techniques achieve a similar performance, since the randomness introduced by the channel cannot be mitigated using only $10$ antenna elements.
@@ V Conclusion
This paper introduced an age-aware edge-blind over-the-air FL framework that is efficient and avoids the need for client-side CSI estimation. The PS aggregates the transmitted updates using multi-antenna MRC and updates only a small subset of coordinates through a two-step technique (AgeTop-$k$): selecting the $r$-largest model magnitudes, and then updating the $k$ most stale of them. By compressing each update into only $k$ selected coordinates that fit into a single OFDM symbol, the proposed method reduces latency while keeping the most useful information. Our convergence analysis has shown that increasing the number of PS antennas improves convergence rate and accuracy. The analysis has also demonstrated a trade-off between compression quality and transmission quality: larger $k$ reduces compression error but increases noise sensitivity, and hence an optimal value must be employed to balance both effects. Theoretical results have been validated by experimental results. Future work include analytically exploring the optimal value of $k$ based on channel conditions and compression error, as well as extending to a fully-blind (CSI-free) system at both the PS and clients.